OpenStudy (anonymous):

Which of the following is equivalent to |x - 2| = 6? x – 2 = 6 x – 2 = -6 x – 2 = 6 or x – 2 = -6 none of the above

OpenStudy (anonymous):

|x| = x OR -x So, |x-2| = x-2 OR -(x-2) = x-2 OR -x+2 And then, |x-2| = 6 is x-2 = 6 OR -x+2 = 6 Multiply both sides by -1 in the second equation: x-2 = 6 OR x-2 = -6

OpenStudy (anonymous):

thanks